Isabella rode her bike at a constant speed. It took her 2 hours to travel 216km with the wind behind her. the return trip was took her 3 hours riding into the wind
thanks,

Answer by richwmiller(17219)

r*t=d
2x=216,
3y=216,
x=b+w,
y=b-w
b bike speed = 90, w wind = 18, bike with wind x = 108, y bike against wind = 72
You didn't ask a question.